OpenStudy (anonymous):

Hey guys i need help on this problem A spring with a spring constant 73.0 n/m sitting vertical on table. The spring is compress 0.210m and a 0.350kg shoe is place on the spring. When the spring is released the shoe shoots into the air. Neglecting air resistance, how high will the shoe go above where it was on the spring?

OpenStudy (anonymous):

F = -kx (from newtons law and hookes law) F = ma ma = -kx a = -kx/m velocity when it returns back to equilibrium (from constant acceleration equations) V^2 = 2a height reached 0 = V^2 + 2a times x solve for x

OpenStudy (anonymous):

the a in the last question is the gravitational constant "g" not the acceleration due to the spring "a", I should have made that clearer

OpenStudy (anonymous):

so I need to replace the "a" with g are gravitational right?

OpenStudy (anonymous):

so your final equation will be 0 = sqrt(-2kx/m) + 2gx i forgot to square root the v^2 for the last equation

OpenStudy (anonymous):

ffs 0 = sqrt (2k times compression) +2gx solve for x

OpenStudy (anonymous):

so the f=-kx are the same as f=ma or two different problem

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

Force is the same regardless whats doing it. its always kg m s^-2 so you can always equate them

OpenStudy (anonymous):

how do i start off the problem im totally lost lol

OpenStudy (lastdaywork):

The question can be solve by two approaches - - You can use energy balance (1 step method) - You can use (vertical) projectile motion (2 step method) Which one would you prefer ??

OpenStudy (anonymous):

projectile motion

OpenStudy (lastdaywork):

Step 1: Find the velocity of the shoe at the point it leaves contact with the spring; or in other words, when the spring returns to its normal length. Can you find it on your own ??

OpenStudy (anonymous):

so what i have is f=73.0 x 0.210m f= 153.3

OpenStudy (anonymous):

f=ma f= 153.3 x 0.350 = 438.0

OpenStudy (lastdaywork):

Have you been taught Work and Energy or SHM ??

OpenStudy (anonymous):

i think work and energy like high school stuff

OpenStudy (lastdaywork):

You have to calculate work done by the spring on the shoe (= change in its KE) You cannot use the equations of motion as acceleration won't be constant.

OpenStudy (anonymous):

so how do I do that can you show me?

OpenStudy (lastdaywork):

Which grade do you study in ??

OpenStudy (anonymous):

junior in highschool

OpenStudy (lastdaywork):

Right now; I can only give you the formula for velocity - v = A*ω where; v = velocity (at the point when the spring is in equilibrium) A (called amplitude) = The distance to which the spring is (initially) compressed or stretched ω (called angular frequency) = √(k/m) k = spring constant m = mass of the object Now find 'v' in this case. PS: You don't really have to worry about the names of the symbols; I gave them for general knowledge :D

OpenStudy (anonymous):

ohh okay thank you

OpenStudy (lastdaywork):

Then, you can use the theory of projectile motion to find the maximum height. "...how high will the shoe go above where it was on the spring? " So, don't forget to add 'A' into the answer you get from projectile.

OpenStudy (anonymous):

thank you I understand now

OpenStudy (lastdaywork):

@Anthony91 Consider closing the question..